OpenStudy (samsan9):

an empty propane tank dropped from a hot air balloon hits the ground with a speed of 148.3 m/s. From what height was the tank released?

OpenStudy (samsan9):

@fateal

OpenStudy (fateal):

ok so this is the formla.... t=d/v distance over velocity... will i think that we need to adjust it tho o.e do u have one for this question?

OpenStudy (fateal):

do u have more info about this question?

OpenStudy (samsan9):

tbh idk one but i am looking through my equation sheet

OpenStudy (fateal):

oh ok...cus that will be easier :)

OpenStudy (samsan9):

h=1/2gt^2

OpenStudy (samsan9):

i know g=9.8

OpenStudy (fateal):

so its going to be h=1/2 (9.8) 148.3 ^ 2 does ^ mean multiply ?

OpenStudy (samsan9):

it means that its squared

OpenStudy (fateal):

oh ok... so did i plug them right?

OpenStudy (samsan9):

hold up let me try it

OpenStudy (fateal):

ok

OpenStudy (fateal):

i got this: 107765.161 so i dont think so??

OpenStudy (foolaroundmath):

Using \(v_f^2 - v_i^2 = 2as \) and taking downward direction to be positive, Assume the tank was released from rest (i.e. \(v_i = 0\)) \( s = \text{height}, \;\;\; v_f = \text{final speed} = 148.3, \; \;\;\;a = g = 9.8\) Plug it in, \(\displaystyle s = \frac{v^2}{2g} = \frac{148.3^2}{2\times9.8} = 1122.08\text{ m}\)

OpenStudy (samsan9):

um ill just put both answers and ask if my friends got it XD

OpenStudy (fateal):

oh ok x3 i was trying to use https://www.mathway.com/ but ok

OpenStudy (samsan9):

lol thank you guys

OpenStudy (fateal):

np x3